# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v2

@@ Abstract
When the quality of naive prompts is carefully optimized by human experts, the task performance of large language models (LLMs) can be significantly improved. However, expert-based prompt optimizations are expensive. Herein, some works have proposed Automatic Prompt Optimization (APO), to optimize naive prompts according to task outputs of given in-box testing models, with the help of advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, existing schemes suffer from poor generalization ability and privacy risk. To this end, we collect the first large-scale Prompt Optimization Preference dataset (POP), fine-tune offline local LLM-based optimizers, then fairly test with various downstream models. Our method allows accurate optimization of the core task instruction part within the naive prompt in a model-agnostic manner, and thus is named Free-from Instruction-oriented Prompt Optimization (FIPO). In specific, FIPO uses a modular APO template that dynamically integrate the naive task instruction, optional instruction responses, and optional ground truth to produce finely optimized prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. Leveraging insights from the data with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of FIPO framework across five public benchmarks and three testing models [^1] Check codes and data here: https://github.com/LuJunru/FIPO_Project..
@@ 2.4 Fine-tuning Strategies
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix A..
@@
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix A and  D..
@@ 3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix B..
@@ 3.1.2 Foundation models
GPT-3.5-turbo Llama2-7B Tulu2-7B
BBH Tasks Naive APE PromptAgent FIPO Naive APE PromptAgent FIPO Naive PromptAgent FIPO GPT-4
Penguins 0.595 0.747 0.797 0.614 0.436 0.402 0.426 0.458 0.336 0.470 0.525 0.542
Geometry 0.227 0.490 0.670 0.738 0.536 0.499 0.544 0.542 0.089 0.350 0.415 0.397
Epistemic 0.452 0.708 0.806 0.865 0.611 0.680 0.775 0.754 0.581 0.691 0.644 0.603
Object Counting 0.612 0.716 0.860 0.660 0.318 0.347 0.312 0.350 0.690 0.577 0.637 0.674
Temporal 0.720 0.856 0.934 0.756 0.362 0.440 0.643 0.683 0.155 0.176 0.210 0.176
Causal Judgment 0.470 0.570 0.670 0.759 0.547 0.583 0.547 0.616 0.832 0.882 0.891 0.888
Average 0.513 0.681 0.790 0.732 0.468 0.492 0.541 0.567 0.447 0.524 0.554 0.547
@@ 3.2.1 Efficacy of FIPO
General improvements of FIPO. We first check if FIPO-optimized prompts have general gains on different downstream generators in Table 2, and conclude that our optimized instruction-oriented prompts are generally beneficial. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average. The improvements reduced as the generator grows larger. This may be due to the fact that larger LLMs has richer intrinsic world knowledge, therefore rely less on the enhanced instructions.
Inference Per Test
Dataset Training Fee Time
APE \$0 \$0 \$5 2h
PromptAgent \$0 \$0 \$5 2h
GPT-4 \$0 \$0 \$4 1h
FIPO \$300 10h \$0 30s
Table 4: Cost report of different methods.
Comparable optimization capability against online ad-hoc APO, even better. We would like to compare the local FIPO optimizer with previous methods and direct prompt optimization using GPT-4. Table 3 reports experimental results on six BBH tasks. When tested with in-box generator GPT-3.5-turbo, PromptAgent obtain significant 28% improvements over all tasks, compared with naive prompts. Although our FIPO optimizer does not use testing feedback from GPT-3.5-turbo, we still maintain over 22% optimization performance. On the other hand, when apply two ad-hoc APO methods on out-of-box Llama2-7B and Tulu2-7B models, we find their enhancement shrinks more than 10%. In contrast, FIPO remains efficacy over different testing generators, and even beats GPT-4.
In terms of cost, shown in Table 4, building FIPO optimizer will have \$300 one-time collecting cost and 10 hours training cost, which equals to 75 usages of GPT-4. However, once FIPO models are prepared, we save a lot of time as the optimization can be done locally. FIPO leads to less overall costs when optimizing over thousands of prompts.
@@ 3.2.2 Fine-tuning Analysis
SFT vs. DPO/IPO vs. IPL. When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task. While when contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%. In terms of different preference fine-tuning methods, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline, which may due to its regularized design in Eq. 10. We analyze more fine-tuning details and the self-rewarding benefits of IPL in Appendix C and  D.
Dataset diversification is necessary. In the bottom of Table 5, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The corrupted optimizer contributes to a decrease on GSM8K and multi-choice benchmarks, while improves the BBH.
@@
In Table 6, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand and apply the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer. We manually confirm that this issue mostly occurs in GSM8K and BBH’s math calculation questions, happening <10%, in Appendix E.
@@ 5 Conclusion
Appendix A Dataset Diversification
Appendix B HyperParameters and Training Cost
@@
Appendix C Entire fine-tuning results
Generation Multi-choice
Naive 24.77 36.21 73.35 51.17 51.22 47.79
SFT 13B 19.91 31.04 71.61 47.68 49.57 44.93
70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO 13B 17.17 30.90 68.50 44.87 46.64 42.69
70B 27.74 35.56 74.17 54.93 52.73 49.07
70B-e1 25.38 33.86 74.39 52.42 52.44 48.12
IPL-DPO 70B-e2 24.28 36.10 73.84 51.22 52.04 48.23
70B-e3 25.13 35.25 74.95 50.46 52.12 48.10
IPO 13B 18.42 33.55 72.03 49.14 48.20 44.93
70B 25.00 39.21 76.84 56.01 54.29 50.94
70B-e1 25.99 34.09 74.66 53.62 52.33 48.30
IPL-IPO 70B-e2 26.70 38.07 76.28 54.38 54.81 50.81
70B-e3 26.67 39.60 77.11 56.71 56.02 52.13
Table 8: Complete evaluation results of different fine-tuning strategies, tested by the same Tulu2-7B model. The “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model).
We present entire fine-tuning results of our various experiments in Table 8. Apart from conclusions reported in Section 3.2.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected. However, all SFT models perform worse than the naive baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe that 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend. We suppose the regularized design in Equation 10 contributes to stable improvements.
@@ Appendix C Entire fine-tuning results
Appendix D Additional details of IPL
@@
As listed in the bottom of Table 8, IPL brings further benefits for IPO training. Compared with basic IPO preference tuning, to pump them into our IPL pipeline can advance to better accuracy, exhibiting a steeper upward growth of 2% per epoch.
we also report the accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in IPL in Table 9. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. We use 5% training data as the validation set, and find 100% classification accuracy. Based on this, the optimizer accurately update naive prompts $\bm{x}^{n}$ with self-generate new prompts $\bm{x}^{n+}$, introduced from line 9 to line 11 in Algorithm 1. Nevertheless, the optimizer model gradually update new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. We observe only 2.4% naive prompts and responses samples are upgraded in the end, while the self-game approach significantly promotes the optimization of downstream test prompts, as mentioned in Table 8.

### v3

@@ Abstract
When the quality of naive prompts is carefully optimized by human experts, the task performance of large language models (LLMs) can be significantly improved. However, expert-based prompt optimizations are expensive. Herein, some works have proposed Automatic Prompt Optimization (APO), to optimize naive prompts according to task outputs of given in-box testing models, with the help of advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, existing schemes suffer from poor generalization ability and privacy risk. To this end, we collect the first large-scale Prompt Optimization Preference dataset (POP), fine-tune offline local LLM-based optimizers, then fairly test with various downstream models. Our method allows accurate optimization of the core task instruction part within the naive prompt in a model-agnostic manner, and thus is named Free-from Instruction-oriented Prompt Optimization (FIPO). In specific, FIPO uses a modular APO template that dynamically integrate the naive task instruction, optional instruction responses, and optional ground truth to produce finely optimized prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. Leveraging insights from the data with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of FIPO framework across five public benchmarks and six testing models [^1] Check codes and data here: https://github.com/LuJunru/FIPO_Project..
@@ 2.4 Fine-tuning Strategies
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix B..
@@
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix B and  C..
@@ 3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix A..
@@ 3.1.2 Foundation models
Figure 4: The FIPO-optimized prompts help various downstream testing LLMs (X-axis) gain more promising improvements, compared with other prompt optimization approaches (shown by the bars). We specifically annotate the improvements of FIPO against naive prompts from the original dataset (). More details: Appendix G and  D.
@@ 3.2.1 Efficacy of FIPO
General improvements of FIPO. It can be concluded that FIPO-optimized prompts have general gains on different downstream generators across five public benchmarks, shown in Table 2. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average.
Comparable optimization capability against online ad-hoc APO, even better. We would like to compare the local FIPO optimizer with previous methods and direct prompt optimization using GPT-4. Figure 4 reports experimental results on six BBH tasks, following the experimental settings in PromptAgent work Wang et al. (2023a). Our FIPO method takes the lead in all downstream tests, except for the in-box tester GPT-3.5-turbo, which is ad-hocly included during the iterative prompt optimization in APE and PromptAgent. In specific, the final average improvements on two open-source 70B models are around 3% to 5%, compared with more than 10% gains on two open-source 7B models. We can notice that as the tested open-soruce model grows larger and stronger, the effectiveness of all prompt optimization methods significantly decreases, which may be due to the firmness of the larger model’s inherent knowledge. As for proprietary GPT-3.5 and GPT-4, we found that prompt optimization seems to be more beneficial to them. Prompts optimized by our FIPO can help GPT3.5 improve the final average effect up to 22%, and maintain an effect of about 8% on GPT4o.
@@ 3.2.2 Fine-tuning Analysis
SFT vs. DPO/IPO vs. IPL. When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task. While when contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%. In terms of different preference fine-tuning methods, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline, which may due to its regularized design in Eq. 10. We analyze more fine-tuning details and the self-rewarding benefits of IPL in Appendix D.
Dataset diversification is necessary. In the bottom of Table 3, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The ablated optimizer weakens all benchmarks, except BBH, which is due to its unique symbolic reasoning pattern (Appendix E).
@@
In Table 4, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand and apply the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer. We manually confirm that this issue mostly occurs in GSM8K and BBH’s math calculation questions, happening <10%, in Appendix H.
@@ 5 Conclusion
Appendix A HyperParameters and Training Cost
@@
Figure 5: An overview of our dataset diversification step. It is recommended to view the details with colors.
Appendix B Dataset Diversification
@@
Generation Multi-choice
Naive 24.77 36.21 73.35 51.17 51.22 47.79
SFT 13B 19.91 31.04 71.61 47.68 49.57 44.93
70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO 13B 17.17 30.90 68.50 44.87 46.64 42.69
70B 27.74 35.56 74.17 54.93 52.73 49.07
70B-e1 25.38 33.86 74.39 52.42 52.44 48.12
IPL-DPO 70B-e2 24.28 36.10 73.84 51.22 52.04 48.23
70B-e3 25.13 35.25 74.95 50.46 52.12 48.10
IPO 13B 18.42 33.55 72.03 49.14 48.20 44.93
70B 25.00 39.21 76.84 56.01 54.29 50.94
70B-e1 25.99 34.09 74.66 53.62 52.33 48.30
IPL-IPO 70B-e2 26.70 38.07 76.28 54.38 54.81 50.81
70B-e3 26.67 39.60 77.11 56.71 56.02 52.13
IPL-IPO-gen 70B-e3 22.72 41.91 74.53 53.60 52.74 50.23
IPL-IPO-partial 70B-e3 23.08 40.03 76.22 54.29 51.99 49.59
Table 6: Complete evaluation results of different fine-tuning strategies, tested by the same Tulu2-7B model. The “-eN” notation refers to the “N”-th round of iteration in IPL training (e.g., 13B-e2 refers to the second iteration of fine-tuning the 13B model). “-gen” and “-partial” refer to two ablation experiments in accordance with Section 3.2.2.
Appendix C Additional details of IPL
@@
we report the accuracy of fine-tuned discrimination ability and proportions of more rigours samples involved in IPL in Table 7. We can notice that the optimizer can easily handle the binary discrimination task with simultaneous training on the additional discrimination data. We use 5% training data as the validation set, and find 100% classification accuracy. Based on this, the optimizer accurately update naive prompts $\bm{x}^{n}$ with self-generate new prompts $\bm{x}^{n+}$, introduced from line 9 to line 11 in Algorithm 1. Nevertheless, the optimizer model gradually update new superior prompt samples for dynamic optimization with a conservative attitude, as shown in the “Selection” column. We observe only 2.4% naive prompts and responses samples are upgraded in the end, while the self-game approach significantly promotes the optimization of downstream test prompts, as mentioned in Table 6.
@@
Appendix D Entire fine-tuning results
We present entire fine-tuning results of our various experiments in Table 6. Apart from conclusions reported in Section 3.2.2, we further summarize several findings: (1) When fine-tuned with SFT approaches, 70B models perform better than 13B models as expected. However, all SFT models perform worse than the naive baseline; (2) In terms of two basic preference optimization training (DPO and IPO), we observe that 70B models are still superior to 13B models, and only the former ones obtains super-human results; (3) When it comes to our IPL approaches, there are differences when combined with DPO or IPO accordingly. IPL-DPO keeps a similar conclusion with DPO. As for IPL-IPO, we particularly find its full training works the best, and obtain an obvious growing trend, exhibiting a steeper upward growth of 2% per epoch. We suppose the regularized design in Equation 10 contributes to stable improvements.
6 BBH Tasks following PromptAgent Work Wang et al. (2023a)
Downstream LLM Optimizer Penguins Geometry Epistemic Object Counting Temporal Causal Judgment Average
Naive 0.436 0.536 0.611 0.318 0.362 0.547 0.468
APE 0.402 0.499 0.680 0.347 0.440 0.583 0.492
Llama2-7B PromptAgent 0.426 0.544 0.775 0.312 0.643 0.547 0.541
Touvron et al. (2023) GPT-4 0.444 0.578 0.640 0.392 0.536 0.598 0.531
FIPO 0.458 0.542 0.754 0.350 0.683 0.616 0.567
Naive 0.336 0.089 0.581 0.690 0.155 0.832 0.447
APE 0.313 0.400 0.627 0.525 0.189 0.882 0.489
Tulu2-7B PromptAgent 0.470 0.350 0.691 0.577 0.176 0.882 0.524
Ivison et al. (2023) GPT-4 0.542 0.397 0.603 0.674 0.176 0.888 0.547
FIPO 0.525 0.415 0.644 0.637 0.210 0.891 0.554
Naive 0.752 0.460 0.720 0.499 0.912 0.651 0.666
APE 0.740 0.533 0.713 0.530 0.897 0.643 0.676
Llama3-70B-Instruct PromptAgent 0.748 0.501 0.667 0.524 0.946 0.687 0.687
Dubey et al. (2024) GPT-4 0.756 0.478 0.726 0.472 0.912 0.644 0.665
FIPO 0.765 0.503 0.678 0.539 0.923 0.716 0.692
Naive 0.593 0.540 0.698 0.602 0.879 0.767 0.680
APE 0.614 0.496 0.718 0.617 0.888 0.790 0.687
Qwen2-72B-Instruct PromptAgent 0.658 0.533 0.717 0.622 0.898 0.779 0.722
Yang et al. (2024) GPT-4 0.630 0.540 0.700 0.642 0.876 0.759 0.691
FIPO 0.617 0.522 0.792 0.591 0.836 0.768 0.731
Naive 0.595 0.227 0.452 0.612 0.720 0.470 0.513
APE 0.747 0.490 0.708 0.716 0.856 0.570 0.681
GPT-3.5-turbo PromptAgent 0.797 0.670 0.806 0.860 0.934 0.670 0.790
(In-box Tester) GPT-4 0.632 0.589 0.840 0.647 0.821 0.550 0.680
FIPO 0.614 0.738 0.865 0.660 0.756 0.759 0.732
Naive 0.859 0.472 0.820 0.680 0.990 0.750 0.762
APE 0.862 0.691 0.844 0.650 0.988 0.749 0.797
GPT-4o PromptAgent 0.855 0.750 0.886 0.733 0.982 0.712 0.820
GPT-4 0.840 0.737 0.925 0.692 0.982 0.700 0.813
FIPO 0.876 0.810 0.891 0.684 0.965 0.840 0.844
Appendix E Specific ablation discussion on BBH
We notice that the BBH dataset behaves significantly different from other datasets in the ablation experiments, as shown by the last lines of Table 6. The corrupted optimizer IPL-IPO-gen and IPL-IPO-partial gain suboptimal performances across all benchmarks, except BBH. We suppose this may be because the BBH contains a large number of symbolic reasoning tasks, and this unique pattern is not found in other testing datasets.
For instance, one of the typical subtask in BBH is “Name a Geometry”, which asks to “name geometric shapes from their SVG paths”. A typical example is listed here: “This SVG path element <path d="M 59.43,52.76 L 75.49,27.45 L 54.92,4.40 M 54.92,4.40 L 23.70,7.77 L 15.15,42.15 L 34.51,57.44 L 59.43,52.76"/> draws a ”. Therefore, we suspect that when further using discrete natural language instructions to diversify our training data, this may affect subtasks in BBH that involve specific symbolic reasoning patterns.
Nevertheless, when examining the six representative subtasks in BBH, FIPO still achieves a clear advantage over other leading prompt optimization methods, as shown by the Table 8.
Appendix F Detailed results of 6 BBH subsets
We provide the performance of various optimizers on different downstream testers, including the evaluation results of 6 BBH tasks in Table 8. Here are some key points of the result analysis:
  • Overall Performance. The FIPO optimizer performs well in most combinations, usually leading in each task and average score. PromptAgent and GPT-4 optimizers perform well on some specific tasks, but are slightly inferior to FIPO in overall average scores.
  • Optimizer Comparison. When comparing different optimizers, FIPO outperforms other methods across multiple downstream LLMs. While PromptAgent occasionally achieves high scores in individual tasks, its average performance is generally lower than FIPO. Similarly, the GPT-4 optimizer shows competitive results in certain scenarios but does not maintain the same level of consistency as FIPO. On the other hand, the Naive prompt and APE optimizer often lag behind in most tasks. For example, on GPT-4o, APE’s average score is 0.797, lower than FIPO’s 0.844.
  • Task-specific Performance. On a task level, FIPO excels in many tasks. For instance, in the "Temporal" and "Causal Judgment" tasks under Llama2-7B, it achieved the highest scores of 0.683 and 0.616, respectively. It also secured the top scores of 0.525 and 0.891 in the "Penguins" and "Causal Judgment" tasks under Tulu2-7B. For GPT-4o, FIPO achieved the highest scores in the "Geometry" and "Causal Judgment" tasks, with 0.810 and 0.840, respectively. Similarly, in the "Epistemic" task under Qwen2-72B-Instruct, FIPO led with a score of 0.792. Overall, FIPO optimizer demonstrates balanced and strong performance across various tasks.
In summary, the FIPO optimizer performed best on a variety of downstream LLMs, especially on GPT-4o and Llama3-70B-Instruct, where its stability and efficiency were remarkable. The PromptAgent and GPT-4 optimizers also performed well on specific tasks, but were not as comprehensive as FIPO overall. The Naive prompt and APE optimizer performed relatively poorly.
Appendix G Cost comparison: FIPO vs. others
In terms of overall cost against other aforementioned optimization approaches, shown in Table 9, building FIPO optimizer will have \$300 one-time collecting cost and 10 hours training cost, which equals to 75 usages of GPT-4. However, once FIPO models are prepared, we save a lot of time as the optimization can be done locally. FIPO leads to less total costs when optimizing over massive prompts.
In specific, we explain the cost number in Table 9 column by column. The first column refers to the cost of constructing data. Since APE, PromptAgent and directly using GPT-4 are all training-free methods, they do not have the cost of constructing training data. Similarly, the second column refers to the time cost of training, and only our local fine-tuned FIPO optimizer covers that. As for the inference cost in the third and fourth columns, we report the inference cost of the entire representative “penguins_in_a_table” subtask in BBH, which includes 149 test samples. In particular, it worth to note that the inference cost of PromptAgent is consistent with its official report [^6] https://github.com/XinyuanWangCS/PromptAgent/. In addition, as the only local optimizer, FIPO also includes a one-time inference deployment cost. However, in our practice, since the time for renting a server and deploying a model does not exceed 3 minutes in total, this part of the cost is basically negligible.
Appendix H Overwhelmed Cheating Notes
Inference Per Test
Dataset Training Fee Time
APE \$0 \$0 \$5 2h
PromptAgent \$0 \$0 \$5 2h
GPT-4 \$0 \$0 \$4 1h
FIPO (IPL-IPO-70B) \$300 \$60 \$0 30s
Table 9: Cost report of different methods.